OpenStudy (anonymous):

-27b^7 divided by 9. can any one hepl me with dividing monomials

OpenStudy (anonymous):

can you more specific (-27b)^7/9 or -27(b^7)/9

OpenStudy (anonymous):

just do -27b multiplied by itself 7 times. than divide it by 9

OpenStudy (anonymous):

first (-27b)^7/9 =-27^7/9 *b^7=(-9*3)^7/9*b^7=-(9^7*3^7)/9*b^7=-(9^6*3^7)*b^7 second -27(b^7)/9 =-3b^7

OpenStudy (anonymous):

im still confused ;-;

OpenStudy (anonymous):

\[\frac{ -27b^7 }{ 9 }\] \[-3b ^{7}\]